• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name C. Keith

    Log entry time 08:02:25 on March 18, 2014

    Entry number 386957

    keyword=Cryotarget fans

    Two attempted cool-downs of the cryotarget occurred on Th and Fr of last week (March 13 and 14). Both cool-downs were aborted due to problems with the fans used to circulate the target fluids in Loop 1 and Loop 2. Loop 3 is not in use for this installation. Both fans were checked and operated normally at room temperature.

    On Thursday, Loop 1 was cooled with H2 in the loop. The standard cool-down procedure utilizing reverse flow of cold 20K return gas thru warm return was ineffective due to some plumbing changes at ESR, so a more direct procedure was undertaken. Cooling then proceeded as normal and the Loop 1 fan was turned on (24 Hz) at approximately 100 K. Current draw seemed high at 4-5 Amps. Based on temperature distributions in the loop, we felt the fan was turning backwards, and so we reversed its direction. At 40 K or so, the fan stopped working because it was drawing more than the current limit of 5 Amps. Attempts to run it with slightly higher current or in the original sense of rotation were unsuccessful. It was felt that perhaps the H2 gas was contaminated, so the cold gas was pumped away as the target warmed.

    On Friday another attempt was made to cool Loop1 with H2. This time we reached about 24 K and condensed approximately 1/3 of the H2 inventory, but again the fan stopped working due to excessive current. We then tried cooling Loop 2 with helium gas (house helium), but once again the fan stopped around 50 K or so.

    We are pursuing two possible explanations: contamination of house helium gas, which was used for pump/purge of both Loop 1 and 2, or thermal contraction problems with both fans. Gas samples from both loops were taken and will be analyzed. Both fans were spun over the weekend at room temperature in hopes of "running in" the (cryogenic) bearings.



    A copy of this log entry has been emailed to: jpchen@jlab.org, smithg@jlab.org